### v2

# YITP-24-61

Abstract
In the absence of gravity, Coleman’s theorem states that the $O(4)$-symmetric instanton solution, which is regular at the origin and exponentially decays at infinity, gives the lowest action. Perturbatively, this implies that any small deformation from $O(4)$-symmetry gives a larger action. In this letter we investigate the possibility of extending this theorem to the situation where the $O(4)$-symmetric instanton is singular, provided that the action is finite. In particular, we show a general form of the potential around the origin, which realizes a singular instanton with finite action. We then discuss a concrete example in which this situation is realized, and analyze non-trivial anisotropic deformations around the solution perturbatively. Intriguingly, in contrast to the case of Coleman’s instantons, we find that there exists a deformed solution that has the same action as the one for the $O(4)$-symmetric solution up to the second order in perturbation. Our result implies that there exist non-$O(4)$-symmetric solutions with finite action beyond Coleman’s instantons, and gives rise to the possibility of the existence of a non-$O(4)$-symmetric instanton with a lower action.
Introduction. The study of instanton solutions in quantum field theory has long been a cornerstone in understanding non-perturbative phenomena. Instantons, as classical solutions to the equations of motion in Euclidean space, offer physical insights into the tunneling process from a metastable (false) vacuum to a stable (true) vacuum [1, 2]. In a model with a single field $\Phi$, one employs the semiclassical approximation to express the decay rate $\Gamma$ per unit time per unit volume ($\mathcal{V}$) as
$\Gamma/\mathcal{V}=A\,e^{-B[\Phi]}\;,$ (1)
where $A$ is the prefactor that contains $\hbar$ corrections, and $B[\Phi]$ is the bounce action defined in terms of the Euclidean action $S[\Phi]$ as
$B[\Phi]\equiv S[\Phi]-S[\Phi_{\rm FV}]\;,$ (2)
with $\Phi_{\rm FV}$ being the location of the false vacuum. Note that in the semiclassical limit one evaluates the action $S[\Phi]$ on a classical solution subjected to appropriate boundary conditions. Indeed, it was mathematically proven in [3] that in the absence of gravity the $O(4)$-symmetric solutions result in the minimal bounce action (the most probable tunneling process) under certain conditions. Those solutions were actually assumed to satisfy the boundary conditions,
$\Phi(\rho)\big{|}_{\rho\rightarrow\infty}=\Phi_{\rm FV}\;,\quad\frac{\mathrm{d}\Phi(\rho)}{\mathrm{d}\rho}\bigg{|}_{\rho\rightarrow 0}=0\;,$ (3)
where $\rho$ denotes the radial coordinate in the 4-dimensional Euclidean space. The second condition in (3) guarantees the regularity of the solution at the center of the true-vacuum bubble. We call Euclidean solutions that satisfy the condition (3) Coleman’s instantons. There have been many studies on the properties on various types of Coleman’s instantons, e.g. [4, 5, 6, 7].
Recently, a class of (unbounded) potentials in an arbitrary spacetime dimension was identified, for which the Coleman instantons do not exist [8, 9, 10]. See [11, 12, 13, 14] for new instantons*1*1*1The authors introduced the UV cutoff to make the action finite. and pseudo bounces which do not satisfy the Coleman boundary conditions, Eq. (3). In this letter we investigate another case where $O(4)$-symmetric Euclidean solutions are singular but their action remains finite.*2*2*2This situation is reminiscent of the Hawking-Turok instanton [15] where gravity is taken into account. In this case, the instanton solution diverges as $\rho\rightarrow 0$, but the action is finite. Namely, we consider $O(4)$-symmetric instantons that do not obey the second condition of Eq. (3), hence they are singular at the true-vacuum bubble. Also, similar to [8, 9], we are interested in an unbounded potential, which violates one of the assumptions for Coleman’s theorem. Moreover, we explicitly show that even if small and regular deformations are taken into account, the total action is still given by that computed on the $O(4)$-symmetric ones. We do not include gravity in this letter.*3*3*3We do not consider the backreaction on the metric due to scalar fluctuations.
Setup. Let us start with a real scalar field model with the Euclidean action,
$S=\int\mathrm{d}^{4}x\sqrt{g}\bigg{[}\frac{1}{2}\,g^{\mu\nu}\partial_{\mu}\Phi(x)\partial_{\nu}\Phi(x)+V(\Phi)\bigg{]}\;,$ (4)
where $V(\Phi)$ is the potential. We consider the 4d flat Euclidean metric,
$\mathrm{d}s^{2}=\mathrm{d}\tau^{2}+\mathrm{d}\vec{x}^{2}=\mathrm{d}\rho^{2}+\rho^{2}(\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\Omega_{2}^{2})\;,$ (5)
where $\tau$ is a Euclidean time, $\rho\equiv\sqrt{\tau^{2}+\vec{x}^{2}}$ and $\mathrm{d}\Omega_{2}^{2}\equiv\mathrm{d}\chi^{2}+\sin^{2}\chi\,\mathrm{d}\phi^{2}$. Notice that although the metric after the second equality in (5) is expressed in the manifestly $O(4)$-symmetric form, the model may admit solutions which are not $O(4)$-symmetric. It may also admit solutions which are singular at $\rho=0$.
From the action (4), the equation of motion (EoM) of $\Phi$ reads
$\Box\Phi-\frac{\mathrm{d}V(\Phi)}{\mathrm{d}\Phi}=0\;,$ (6)
where $\Box\equiv g^{\mu\nu}\partial_{\mu}\partial_{\nu}$. The EoM above in general describes dynamics of instanton solution with prescribed boundary conditions: one at the origin of the true-vacuum bubble ($\rho=0$) and the other at false vacuum ($\rho\rightarrow\infty$).
In order to construct the potential $V(\Phi)$, we first consider the $O(4)$-symmetric instanton: $\Phi=\bar{\Phi}(\rho)$. Eq. (6) then reduces to
$\bar{\Phi}^{\prime\prime}+\frac{3}{\rho}\bar{\Phi}^{\prime}-\frac{\mathrm{d}V}{\mathrm{d}\Phi}\bigg{|}_{\Phi=\bar{\Phi}}=0\;,$ (7)
where ${}^{\prime}\equiv{\rm d}/{\rm d}\rho$. The corresponding action reads
$S_{0}=2\pi^{2}\int_{0}^{\infty}\mathrm{d}\rho~{}\rho^{3}~{}\bigg{[}\frac{1}{2}\bar{\Phi}^{\prime 2}+V(\bar{\Phi})\bigg{]}\;,$ (8)
where the factor $2\pi^{2}$ was obtained from the integrals over the angular variables.
Here we allow an instanton to be singular at the origin, or equivalently the true vacuum is located infinitely far away in field space, provided that the action remains finite.*4*4*4Note that there could be a cuspy solution at the origin with the first derivative of the scalar field remains finite. From (8), we see that if $\bar{\Phi}^{\prime 2}(\rho)$ blows up faster than $1/\rho^{3}$ when $\rho\rightarrow 0$, the action will inevitably diverge. Therefore, the finiteness of the action yields the conditions on the behaviors of $\bar{\Phi}^{\prime}$ and $\bar{\Phi}$ at the origin.
From the above reasoning, we find that the general form of the potential as $\rho\rightarrow 0$ with finite action is
$V(\Phi)=-\Lambda^{4}\exp\bigg{(}\frac{2\Phi}{\Phi_{\star}}\bigg{)}\;,$ (9)
where $\Lambda$ and $\Phi_{\star}$ are arbitrary constants. One can straightforwardly verify that Eq. (9) leads to logarithmic divergence of $\bar{\Phi}$ at the origin, but it gives finite action. Indeed, the form (9) violates one of the assumptions of the theorem [3], so that it does not apply in our case. Notice that the potential (9) has not been considered in the literature, and here we do not need to introduce a UV cutoff since our integral over $\rho$ is finite. See [16] for discussion of other types of singularities of $\bar{\Phi}$. Note also that it is straightforward to show that the potential (9) is generic in $d>2$ dimensions.
Small anisotropic deformation. Let us analyze a small anisotropic deformation $\delta\Phi$ of the $O(4)$-symmetric instanton,
$\Phi(x^{\mu})=\bar{\Phi}(\rho)+\epsilon\,\delta\Phi(\rho,\vec{\theta}\,)\;,$ (10)
where $\epsilon$ is the anisotropic parameter satisfying $\epsilon\ll 1$. Notice that $\delta\Phi$ can generally depend on both $\rho$ and the angular variables, $\vec{\theta}=\{\theta,\chi,\phi\}$. The solution $\delta\Phi(\rho,\vec{\theta}\,)$ can be decomposed as
$\delta\Phi(\rho,\vec{\theta}\,)=\sum_{L,M}A_{LM}(\rho)\,Y_{L}^{M}(\vec{\theta}\,)\;,$ (11)
where the function $Y_{L}^{M}(\vec{\theta}\,)$ is the 3-dimensional spherical harmonics [17], obeying
$\Delta_{S^{3}}Y_{L}^{M}(\vec{\theta}\,)=-L(L+2)Y^{M}_{L}(\vec{\theta}\,)\;,$ (12)
with $\Delta_{S^{3}}$ the Laplacian of a 3-sphere and $M$ a multi-index characterizing the magnetic quantum numbers $\{m_{\phi},m_{\chi},m_{\theta}\}$ satisfying $|m_{\phi}|\leq m_{\chi}\leq L\equiv m_{\theta}$.
The radial component $A_{L}(\rho)$ then satisfies
$A^{\prime\prime}_{L}+\frac{3}{\rho}A^{\prime}_{L}-\bigg{[}\frac{\,\mathrm{d}^{2}V}{\,\mathrm{d}\Phi^{2}}\bigg{|}_{\bar{\Phi}}+\frac{L(L+2)}{\rho^{2}}\bigg{]}A_{L}=0\;,$ (13)
where we have omitted the index $M$. This equation determines the dynamics of fluctuations $A_{L}$ on the background $\bar{\Phi}$. Notice that the modes with $L=0$ and $L=1$ are not physical, see [16] for detailed discussion. Straightforwardly, inserting (10) into (4), one arrives at the action up to $\mathcal{O}(\epsilon^{2})$,
$S=S_{0}[\bar{\Phi}]+\epsilon^{2}S_{2}[\delta\Phi]\;,$ (14)
where $S_{0}$ is given by Eq. (8) and
$S_{2}=\frac{1}{2}\int\mathrm{d}^{4}x\sqrt{g}\,\bigg{[}(\partial\delta\Phi)^{2}+\frac{\,\mathrm{d}^{2}V}{\,\mathrm{d}\Phi^{2}}\bigg{|}_{\bar{\Phi}}\delta\Phi^{2}\bigg{]}\;.$ (15)
Note that we have performed the integration by parts and used Eq. (7) to obtain (14). (The action $S_{2}$ is the leading-order correction to the $O(4)$-symmetric on-shell action $S_{0}$.) Note also that the effective mass for $A_{L}$, i.e., $\,\mathrm{d}^{2}V/\,\mathrm{d}\Phi^{2}$ evaluated on $\bar{\Phi}$ is only a function of $\rho$; the only non-trivial angular dependence appears through the fluctuation $\delta\Phi(\rho,\vec{\theta}\,)$. Hence, using the solution (11) in the action $S_{2}$ gives
$S_{2}=\sum_{L\geq 2}\int\mathrm{d}\rho\,\mathcal{L}_{2L}\;,$ (16)
where
$\mathcal{L}_{2L}\equiv\frac{\rho^{3}}{2}\bigg{\{}A_{L}^{\prime 2}+\bigg{[}\frac{L(L+2)}{\rho^{2}}+\frac{\,\mathrm{d}^{2}V}{\,\mathrm{d}\Phi^{2}}\bigg{|}_{\bar{\Phi}}\bigg{]}A_{L}^{2}\bigg{\}}\;,$ (17)
and we have used the orthogonality condition for spherical harmonics. For regular $A_{L}$, the action (16) can be integrated as
$S_{2}=\pi^{2}\sum_{L\geq 2}\rho^{3}A_{L}(\rho)A^{\prime}_{L}(\rho)\bigg{|}_{\rho=0}^{\rho=\infty}\;,$ (18)
where we have performed an integration by parts and used Eq. (13). Notice that the above expression does not hold for the singular $A_{L}$.*5*5*5We discuss this point in our separate work [16]. Note that in the singular case, one may introduce the cutoff to regularize the action as $\rho\rightarrow 0$, see e.g. [13]. We see that the action $S_{2}$ is always finite as long as $A_{L}$ is regular everywhere. Moreover, it indicates that $S_{2}$ is solely determined by the behaviors of $A_{L}$ at the origin and infinity. As mentioned above, our model is beyond Coleman theorem [3] since the $O(4)$-symmetric solution $\bar{\Phi}$ is assumed to be singular at $\rho=0$ and the potential is not admissible everywhere. It may be interesting to have a rigorous proof of the least action with a singular instanton, we leave this to future work. In the following, we present a concrete example which explicitly realizes the case we are interested in.
Example: Piecewise quadratic potential. We study an example, where the potential is
$\frac{V(\Phi)}{\alpha^{4}}=\left\{\begin{matrix}-\Phi_{\star}^{4}\exp(2\Phi/\Phi_{\star})\;,&\Phi\geq 0\\ V_{1}(\Phi)\;,&\Phi_{2}\leq\Phi\leq 0\\ V_{2}(\Phi)\;,&\Phi\leq\Phi_{2}\end{matrix}\right.\;,$ (19)
where
$V_{1}(\Phi)$ $\equiv-\frac{1}{2}m_{1}^{2}(\Phi-\Phi_{\rm P})^{2}-\Lambda_{1}^{4}\;,$ (20)
$V_{2}(\Phi)$ $\equiv\frac{1}{2}m_{2}^{2}(\Phi-\Phi_{\rm M})^{2}-\Lambda_{2}^{4}\;,$ (21)
where $\Lambda$, $m_{1}$, $m_{2}$, $\Phi_{\rm P}$, $\Phi_{\rm M}$, $\Lambda_{1}$, $\Lambda_{2}$, $\Phi_{\star}$ are parameters with mass dimension, $\alpha$ is an overall dimensionless parameter. Notice that $\Phi_{\rm P}$ and $\Phi_{\rm M}$ denote the locations of the maximum and the minimum (false vacuum) in $V_{1}$ and $V_{2}$, respectively. Then, imposing the continuity conditions for the potential and its first derivative at $\Phi=0$ and $\Phi=\Phi_{2}$, one can fix the parameters $\{\Lambda_{1},\Lambda_{2},\Phi_{\rm P},\Phi_{2}\}$ in terms of $\{m_{1},m_{2},\Phi_{\rm M}\}$ such that
$\Phi_{\rm P}$ $=-\frac{2\Phi_{\star}^{3}}{m_{1}^{2}}\;,$ (22)
$\Phi_{2}$ $=-\frac{2\Phi_{\star}^{3}}{m_{1}^{2}+m_{2}^{2}}\bigg{(}1-\frac{m_{2}^{2}\Phi_{\rm M}}{2\Phi_{\star}^{3}}\bigg{)}\;,$
$\Lambda_{1}$ $=\Phi_{\star}\bigg{(}1-\frac{\Phi_{\star}^{2}}{m_{1}^{2}}\bigg{)}\;,$
$\Lambda_{2}$ $=-\frac{2\Phi_{\star}^{3}}{(m_{1}^{2}+m_{2}^{2})}\bigg{[}1-\bigg{(}1+\frac{2\Phi_{\rm M}}{\Phi_{\star}}\bigg{)}\frac{m_{2}^{2}}{2\Phi_{\star}^{2}}$
$\hskip 11.95013pt-\bigg{(}1+\frac{m_{2}^{2}\Phi_{\rm M}^{2}}{2\Phi_{\star}^{4}}\bigg{)}\frac{m_{1}^{2}}{2\Phi_{\star}^{2}}\bigg{]}\;.$
Moreover, the location of the false vacuum must fall into the regime of $V_{2}$, which implies that $\Phi_{\rm M}<\Phi_{2}<0$. This gives the condition*6*6*6One can straightforwardly verify that the condition (23) automatically leads to $\Phi_{2}\leq\Phi_{\rm P}<0$.,
$2+\frac{m_{1}^{2}\Phi_{\rm M}}{\Phi_{\star}^{3}}<0\;.$ (23)
The above condition must be satisfied for any value of $m_{1}$ and $\Phi_{\rm M}$. Notice that in our setup we have $\Phi_{\rm M}<0$. For concreteness, in Fig. 1 we plot the potential (19) with $m_{1}=0.98\Phi_{\star}$, $m_{2}=7.11\Phi_{\star}$, $\Phi_{\rm M}=-2.35\Phi_{\star}$ and $\alpha=0.5$. Note that these values of parameters are chosen such that all the matching conditions of the background solutions are satisfied. In the plot, the red solid line corresponds to the exponential potential, whereas the blue dashed and the green dot-dashed lines refer to the potentials $V_{1}$ and $V_{2}$, respectively. The two black dots represent the matching points at $\Phi=0$ and $\Phi=\Phi_{2}=-2.34\Phi_{\star}$.
Figure 1:  $V(\Phi)/(\alpha\Phi_{\star})^{4}$ in (19) as a function of $\Phi$ with $m_{1}=0.98\Phi_{\star}$, $m_{2}=7.11\Phi_{\star}$ and $\Phi_{\rm M}=-2.35\Phi_{\star}$. The solid red line represents the exponential potential, while the blue dashed and the green dot-dashed lines refer to the potentials $V_{1}$ and $V_{2}$ in Eq. (19) respectively. The two black points in the plot refer to the two matching locations at $\Phi=0$ and at $\Phi=\Phi_{2}=-2.34\Phi_{\star}$. The false vacuum for this particular choice of parameters is located at $\Phi_{\rm M}=-2.35\Phi_{\star}$.
At the background level, in each regime of the potential (19) the solutions $\bar{\Phi}$ can be obtained analytically with their integration constants. Throughout this example, we do not use the thin-wall approximation. In the exponential-potential regime, solving Eq. (7) we find that
$\bar{\Phi}(\tilde{\rho})=-\Phi_{\star}\log(\tilde{\rho})\,;~{}0<\tilde{\rho}\leqslant 1\,,$ (24)
where $\tilde{\rho}\equiv\alpha^{2}\Phi_{\star}\rho$. Notice that in (24) we focus on the singular behavior, so that the regular behavior is disregarded.*7*7*7Of course, once the condition $\Phi^{\prime}_{0}\rightarrow\infty$ at $\rho=0$ is imposed, $\bar{\Phi}$ is uniquely determined by Eq. (24). In the $V_{1}$ regime the solution reads
$\bar{\Phi}(\tilde{\rho})=$ $-\frac{2\Phi_{\star}^{3}}{m_{1}^{2}}+\frac{c_{1}}{\tilde{\rho}}J_{1}\big{(}\frac{m_{1}}{\Phi_{\star}}\tilde{\rho}\big{)}+\frac{c_{2}}{\tilde{\rho}}Y_{1}\big{(}\frac{m_{1}}{\Phi_{\star}}\tilde{\rho}\big{)}\,;$
$1\leqslant\tilde{\rho}\leqslant\tilde{\rho}_{2}\,,$ (25)
where $J_{1}$ and $Y_{1}$ are Bessel functions of the first and second kinds respectively, and $c_{1}$, $c_{2}$ are constants. In the $V_{2}$ regime we have
$\bar{\Phi}(\tilde{\rho})=\Phi_{\rm M}+\frac{c_{3}}{\tilde{\rho}}K_{1}\big{(}\frac{m_{2}}{\Phi_{\star}}\tilde{\rho}\big{)}+\frac{c_{4}}{\tilde{\rho}}I_{1}\big{(}\frac{m_{2}}{\Phi_{\star}}\tilde{\rho}\big{)}\,;~{}\tilde{\rho}\geqslant\tilde{\rho}_{2}\,,$ (26)
where $K_{1}$ and $I_{1}$ are modified Bessel functions of the first and second kinds respectively*8*8*8The Bessel functions $K_{1}(z)$, $I_{1}(z)$ behaves as $e^{-z}$ and $e^{z}$ as $z\rightarrow\infty$., and $c_{3}$, $c_{4}$ are constants. We see that in order for $\bar{\Phi}$ asymptotically goes to $\Phi_{\rm M}$ as $\tilde{\rho}\rightarrow\infty$, we impose $c_{4}=0$. Then, using the continuity conditions at $\tilde{\rho}=1$ ($\bar{\Phi}=0$) the constants $c_{1}$ and $c_{2}$ are fixed in terms of $m_{1}$. At the second matching point $\tilde{\rho}=\tilde{\rho}_{2}$ ($\bar{\Phi}=\Phi_{2}$), one uses the continuities of both $\bar{\Phi}$ and $\bar{\Phi}^{\prime}$ to obtain
$\bigg{(}2+\frac{m_{1}^{2}\Phi_{\rm M}}{\Phi_{\star}^{3}}\bigg{)}H_{1}+H_{2}=0\;,$ (27)
where $H_{1}$ and $H_{2}$ are functions of $m_{1}$, $m_{2}$, and $\tilde{\rho}_{2}$. Hence, $\Phi_{\rm M}$ is also fixed by these parameters. Therefore, inserting the expression for $\Phi_{\rm M}$ into (23), we obtain the inequality,
$H_{3}\left(\frac{m_{1}}{\Phi_{\star}},\frac{m_{2}}{\Phi_{\star}},\tilde{\rho}_{2}\right)\leq 0\;.$ (28)
Here we omit the full expressions of $H_{1}$, $H_{2}$ and $H_{3}$ due to their length. In addition to (28), we identify the solution (26) at $\tilde{\rho}=\tilde{\rho}_{2}$ with $\Phi_{2}$ in (22), i.e.,
$\bar{\Phi}(\tilde{\rho}_{2})=\Phi_{2}\,.$ (29)
We see that for any values of $\tilde{\rho}_{2}>1$ the above conditions must be satisfied in order for the background solution in all regimes can be smoothly connected and asymptotically goes to the false vacuum $\Phi_{\rm M}$. For illustrative purposes, in Fig. 2 we plot the parameter space of $m_{1}/\Phi_{\star}$ and $m_{2}/\Phi_{\star}$ with $\tilde{\rho}_{2}=5$ and $\alpha=0.5$. The yellow region represents the allowed region, obtained by the condition (28) and the green curve refers to the condition $\bar{\Phi}=\Phi_{2}$ at $\tilde{\rho}=\tilde{\rho}_{2}$. After fixing $\tilde{\rho}_{2}$ and imposing the above condition our free parameters are $m_{1}$ and $\alpha$.
Figure 2:  Parameter space of $m_{1}/\Phi_{\star}$ and $m_{2}/\Phi_{\star}$. The condition (28) is satisfied in the yellow region. The green line refers to the condition $\bar{\Phi}(\tilde{\rho}_{2})=\Phi_{2}$, the red curve corresponds to the condition $c_{8}=0$. We choose $\tilde{\rho}_{2}=5$ and $\alpha=0.5$. The black dot is $\{0.98,7.11\}$, for which the fluctuation is regular everywhere (their corresponding $\Phi_{2}$ and $\Phi_{\rm M}$ are respectively given by $\Phi_{2}=-2.34\Phi_{\star},\Phi_{\rm M}=-2.35\Phi_{\star}$). The small figure explicitly shows the intersection point.
Having determined the background O(4)-symmetric solution $\bar{\Phi}$, now we turn to analyze the small deformation $\delta\Phi(\rho,\vec{\theta}\,)$ in (10). In order to solve its radical components $A_{L}$, we introduce a dimensionless variable $f_{L}\equiv\tilde{\rho}^{3/2}A_{L}/\Phi_{\star}$. In the exponential regime of potential (19), solving (13) gives
$f_{L}(\tilde{\rho})=a_{1}\tilde{\rho}^{\frac{1}{2}+\sqrt{j}}+a_{2}\tilde{\rho}^{\frac{1}{2}-\sqrt{j}}\,;~{}0<\tilde{\rho}\leqslant 1\,,$ (30)
where $a_{1}$, $a_{2}$ are constants and $j\equiv(L-1)(L+3)$. Here we focus on the regular behavior at the origin, i.e., $a_{2}=0$, in order for the action to be finite. Note that this choice excludes the possibility that the solution behaves as $1/\sqrt{\tilde{\rho}}$.*9*9*9This is the only singular behavior giving rise to finite action, see Eq. (17). In the $V_{1}$ regime we obtain
$f_{L}(\tilde{\rho})=$ $\sqrt{\tilde{\rho}}\left[c_{5}J_{L+1}\big{(}\frac{m_{1}}{\alpha^{2}\Phi_{\star}}\tilde{\rho}\big{)}+c_{6}Y_{L+1}\big{(}\frac{m_{1}}{\alpha^{2}\Phi_{\star}}\tilde{\rho}\big{)}\right]\,;$
$1\leqslant\tilde{\rho}\leqslant\tilde{\rho}_{2}\;,$ (31)
where $c_{5}$ and $c_{6}$ are constants. Similarly, the solution in the $V_{2}$ regime reads
$f_{L}(\tilde{\rho})=$ $\sqrt{\tilde{\rho}}\left[c_{7}K_{L+1}\big{(}\frac{m_{2}}{\alpha^{2}\Phi_{\star}}\tilde{\rho}\big{)}+c_{8}I_{L+1}\big{(}\frac{m_{2}}{\alpha^{2}\Phi_{\star}}\tilde{\rho}\big{)}\right]\,;$
$\tilde{\rho}\geqslant\tilde{\rho}_{2}\,,$ (32)
where $c_{7}$ and $c_{8}$ are constants. Using the matching conditions for $f_{L}(\tilde{\rho})$ and its first derivative at $\tilde{\rho}=1$ one can fix the constants $c_{5}$ and $c_{6}$ in terms of $a_{1}$, $m_{1}$ and $\alpha$. Furthermore, imposing the matching conditions at $\tilde{\rho}=\tilde{\rho}_{2}$ one can determine $c_{7}$ and $c_{8}$ in terms of $\alpha$, $m_{1}$, $m_{2}$, $\tilde{\rho}_{2}$ and $a_{1}$. Finally, we require the solution (Beyond Coleman’s Instantons) to be regular as $\tilde{\rho}\rightarrow\infty$, this fixes $c_{8}(m_{1},m_{2},\tilde{\rho}_{2},a_{1},\alpha)=0$. For given values of $\tilde{\rho}_{2}$, $\alpha$ and $a_{1}$, this condition provides a relation between $m_{1}$ and $m_{2}$.
Figure 3:  The solutions for $\bar{\Phi}$ (solid line) and $f_{2}(\tilde{\rho})$ (dashed line). The red, blue and green colors refer to the solutions in the exponential potential, $V_{1}$ and $V_{2}$ regimes, respectively. We set $\alpha=0.5$, $m_{1}=0.98\Phi_{\star}$ and $m_{2}=7.11\Phi_{\star}$, so that $\Phi_{2}=-2.34\Phi_{\star}$ and $\Phi_{\rm M}=-2.35\Phi_{\star}$. The two matching points (black dots) are located at $\tilde{\rho}=1$ and $\tilde{\rho}_{2}=5$.
In Fig. 2 we plot the regular condition $c_{8}=0$ with $\tilde{\rho}_{2}=5$ and $\alpha=0.5$, together with the conditions (28) and $\bar{\Phi}(\tilde{\rho})=\Phi_{2}$. In the plot the red line corresponds to the vanishing of the coefficient $c_{8}$ of the solution (Beyond Coleman’s Instantons). The intersection point $\{0.98,7.11\}$ represents the location where all of the conditions [$c_{8}=0$, $\bar{\Phi}(\tilde{\rho})=\Phi_{2}$ and Eq. (28)] are simultaneously satisfied. In other words, for particular values of $\tilde{\rho}_{2}$, $a_{1}$ and $\alpha$ there exists a point where the background $\bar{\Phi}$ and the regular deformation can be realized. Therefore, our independent parameters are $\tilde{\rho}_{2}$ and $\alpha$. For other values of $\tilde{\rho}_{2}$ and $\alpha$ one needs to check the existence of the solutions case by case.
For completeness, in Fig. 3 we plot both the solutions $\bar{\Phi}$ (solid line) and $f_{2}$ (dashed line) as a function of $\tilde{\rho}$ with $\alpha=0.5$, $a_{1}=1$, $m_{1}=0.98\Phi_{\star}$ and $m_{2}=7.11\Phi_{\star}$ such that we have $\Phi_{2}=-2.34\Phi_{\star},\Phi_{\rm M}=-2.35\Phi_{\star}$. Note that in the plot we appropriately rescale $f_{2}$ with the parameter $a_{1}$. The solutions in the exponential potential, $V_{1}$ and $V_{2}$ regimes are labeled by red, blue and green respectively. The black dot denotes the matching points at $\tilde{\rho}=1$ and $\tilde{\rho}=5$. We see that $\bar{\Phi}$ is singular as $\tilde{\rho}\to 0$, while it goes to $\Phi_{\rm M}$ as $\rho\to\infty$. Moreover, the solution $f_{2}(\tilde{\rho})$ is regular everywhere, especially at the origin and infinity. Therefore, in this model there exists the regular solution for small deformations around the singular instanton.
More importantly, the deformation is regular everywhere. Inserting the solution given by (30) and (Beyond Coleman’s Instantons) with $a_{2}=c_{8}=0$ into (18), we find $S_{2}=0$. Therefore, the total action (14) is the same as the one evaluated for the $O(4)$-symmetric instanton. In other words, a small and regular deformation can be regarded as a zero mode on the $O(4)$-symmetric solutions. Our result implies that there exist non-$O(4)$-symmetric solutions with finite action beyond Coleman’s instantons.
Discussion. We studied a possibility of extending the theorem [3] to the case where the $O(4)$-symmetric instanton solution exhibits a singular behavior at $\rho=0$. In particular, we considered the case in which the solution does not satisfy Coleman’s condition (3). We further investigated a small anisotropic deformation around the singular $O(4)$-symmetric instanton. We proposed a concrete example with a piecewise potential given by (19). We then analyzed the solutions for $\bar{\Phi}$, that is singular at $\rho=0$ and goes to false vacuum as $\rho\rightarrow\infty$. Finally, we studied the dynamics of small anisotropic deformation and found that there exists a small and regular deformation around the singular $O(4)$-symmetric solution, which does not contribute to the total bounce action. Our result implies that there exist non-$O(4)$-symmetric solutions with finite action beyond Coleman’s instantons, and gives rise to the possibility of the existence of a non-$O(4)$-symmetric instanton with a lower action. This opens several possibilities to study other cases, where the Coleman’s theorem may not apply. It would be interesting to explore whether the so-called pseudo bounces [10, 14] give lower or higher action than the one in our case, and study their implications.*10*10*10It could be that some off-shell configurations lead to lower actions, but they may correspond to subcritical bubbles. We leave this investigation to future work. Moreover, it is of interest to extend our analysis to the case with gravity. In connection with this, see e.g. [18, 19, 20, 21, 22, 23]. In our separate work [16] we will discuss other types of singular behaviors in addition to logarithmic divergence, and the generic form of the instanton potential.
Acknowledgements. We thank S. Mukohyama, N. Oshita, R. Saito and K. Takahashi for useful discussions. We thank J. R. Espinosa for pointing out an error in the first version of the paper. This work is supported in part by JSPS KAKENHI Nos. JP20H05853 and JP24K00624. M.S. and V.Y. are supported by World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan. Y.Z. is supported by the Fundamental Research Funds for the Central Universities, and by the Project 12475060 and 12047503 supported by NSFC, Project 24ZR1472400 sponsored by Natural Science Foundation of Shanghai, and also by China Scholarship Council.